If possible use diagrams to aid calculations.


4. In a raffle there is one first prize of $100, one second prize of $50 and one third prizeof $10. There are 100 tickets sold and first prize is drawn first, then second, then.

third, without replacement. Find the probability that a person buying one ticket in.

the raffle wins.

(a) first prize;.

(b) a prize;.

(c) at least $50.


1
Expert's answer
2021-06-04T05:44:49-0400

(a) "P=\frac{1}{100}=0.01."

(b) "P=\frac{3}{100}=0.03."

(c) "P=\frac{2}{100}=\frac{1}{50}=0.02."